# Marco Randoa1, Martin Jamesb, Alessandro Verria, Lorenzo Rosascoa, Agnese Seminarab1

### v1

@@ Abstract
We consider the problem of olfactory searches in a turbulent environment. We focus on agents that respond solely to odor stimuli, with no access to spatial perception nor prior information about the odor location. We ask whether navigation strategies to a target can be learned robustly within a sequential decision making framework. We develop a reinforcement learning algorithm using a small set of interpretable olfactory states and train it with realistic turbulent odor cues. By introducing a temporal memory, we demonstrate that two salient features of odor traces, discretized in few olfactory states, are sufficient to learn navigation in a realistic odor plume. Performance is dictated by the sparse nature of turbulent plumes. An optimal memory exists which ignores blanks within the plume and activates a recovery strategy outside the plume. We obtain the best performance by letting agents learn their recovery strategy and show that it is mostly casting cross wind, similar to behavior observed in flying insects. The optimal strategy is robust to substantial changes in the odor plumes, suggesting minor parameter tuning may be sufficient to adapt to different environments.
aMalGa - DIBRIS, University of Genova, ItalybMalGa - DICCA, University of Genova, ItalyThe authors declare no conflict of interest.1To whom correspondence should be addressed. E-mail: marco.rando@edu.unige.it, agnese.seminara@unige.itAuthor’s contributions. M.R., A.V., L.R. and A.S. designed research; M.R., M.J., A.V., L.R. and A.S. performed research; M.R., L.R. and A.S. wrote manuscript.Keywords: Navigation |||| Reinforcement learning |||| Olfaction |||| Turbulence |||| Memory
@@
Bacterial cells localize the source of an attractive chemical even if they hold no spatial perception. They respond solely to temporal changes in chemical concentration and the result of their response is that they move toward attractive stimuli by climbing concentration gradients [1]. Larger organisms also routinely sense chemicals in their environment to localize or escape targets, but cannot follow chemical gradients since turbulence breaks odors into sparse pockets and gradients lose information [2, 3, 4, 5, 6]. The question of which features of turbulent odor traces are used by animals for navigation is natural, but not well understood. Beyond olfaction, some animals could use also prior spatial information to navigate  [7, 8, 9, 10], but if and how chemosensation and spatial perception are coupled is still not clear.
@@
An algorithmic perspective to olfactory navigation in turbulence can shed light on some of these questions. Without aiming at an exhaustive taxonomy, see e.g. [11] for a recent review, we recall some approaches relevant to put our contribution in context. One class of methods are biomimetic algorithms, where explicit navigation rules are crafted taking inspiration from animal behavior. An advantage of these methods is interpretability, in the sense that they provide insights into features that effectively achieve turbulent navigation, for example: odor presence/absence [12, 13, 14, 5]; odor slope at onset of detection [15]; number of detections in a given interval of time [16] and the time of odor onset [17]. On the flip side, in biomimetic algorithms behaviors are hardwired and typically reactive, not relying on any optimality criterion.
@@
A way to tackle this shortcoming is to cast olfactory navigation within a sequential decision making framework [18]. In this context, navigation is formalized as a task with a reward for success; by maximizing reward, optimal strategies can be sought to efficiently reach the target. A byproduct is that most algorithmic choices can often be done in a principled way. Within this framework, some approaches make explicit use of spatial information. Bayesian algorithms use a spatial map to guess the target location and use odor to refine this guess or “belief”. A prominent algorithm for olfactory navigation based on the concept of belief is the information-seeking algorithm [3] akin to exploration heuristics widely used in robotics [19, 20] (see e.g. [21, 22]). Using Bayesian sequential decision making and the notion of beliefs, navigation can be formalized as a Partially Observable Markov Decision Process (POMDP) [23, 24, 25], that can be approximatively solved [26, 27, 28]. POMDP approaches are appealing since beliefs are a sufficient statistics for the entire history of odor detections. However, they are computationally cumbersome. Further, they leave the question open of whether navigation as sequential decision making can be performed using solely olfactory information.
@@
Recently, two algorithms studied navigation as a response to olfactory input alone [29, 30]. In [29] artificial neural networks were shown to learn near optimal strategies, but they were trained on odor cues with limited sparsity, and training with sparse odor cues typical of turbulence remains to be tested. In [30] an approach based on finite state controllers was proposed. Here, optimization was done using a model-based technique, relying on prior knowledge of the likelihood to detect the odor in space, hence still using spatial information. A different model free optimization could also be considered avoiding spatial information but this latter approach also remains to be tested. More generally, all the above approaches manipulate internally the previous history (memory) of odor detections. In this sense they are less interpretable, since the features of odor traces that drive navigation do not emerge explicitly.
Figure 1: Learning a stimulus-response strategy for turbulent navigation. (A) Representation of the search problem with turbulent odor cues obtained from Direct Numerical Simulations of fluid turbulence (grey scale, odor snapshot from the simulations). The discrete position $s$ is hidden; the odor concentration $z_{T}=z(s(t^{\prime}),t^{\prime})|t-T\leq t^{\prime}\leq t$ is observed along the trajectory $s(t^{\prime})$, where $T$ is the sensing memory. (B) Odor traces from direct numerical simulations at different (fixed) points within the plume. Odor is noisy and sparse, information about the source is hidden in the temporal dynamics. (C) Contour maps of olfactory states with nearly infinite memory ($T=2598$): on average olfactory states map to different locations within the plume and the void state is outside the plume. (D) Performance of stimulus-response strategies obtained during training, averaged over 500 episodes. We train using realistic turbulent data with memory $T=20$ and backtracking recovery.
@@
In this paper, we propose a reinforcement learning (RL) approach to navigation in turbulence based on a set of interpretable olfactory features, with no spatial information, and highlight the role played by memory within this context. More precisely, we learn optimal strategies from data by training tabular Q learning [18] with realistic odor cues obtained from state-of-the-art Direct Numerical Simulations of turbulence. From the odor cues, we define features as moving averages of odor intensity and sparsity: the moving window is the temporal memory and naturally connects to the physics of turbulent odors. States are then obtained discretizing such features. Due to sparsity, agents may detect no odor within the moving window. We show there is an optimal memory minimizing the occurrence of this “void state”. The optimal memory scales with the blank time dictated by turbulence as it emerges from a trade off requiring that: (i) short blanks– typical of turbulent plumes– are ignored by responding to detections further in the past, and (ii) long blanks promptly trigger a recovery strategy to make contact with the plume again. We leverage these observations to tune the memory adaptively, by setting it equal to the previous blank experienced along an agent’s path. With this choice, the algorithm tests successfully in distinct environments, suggesting that tuning can be made robustly to enable generalization. The agent learns to surge upwind in most non-void states and to recover by casting crosswind in the absence of detections. Optimal agents limit encounters with the void state to a narrow band right at the edge of the plume. This suggests that the temporal odor features we considered effectively predict when the agent is exiting the plume and point to an intimate connection between temporal predictions and spatial navigation.
Significance
Finding mates or food in the presence of turbulence is challenging because odors constantly switch on and off unpredictably. As a result, it is unclear whether animals couple odor to other sources of information, what are the relevant features of odor stimuli and how they change according to the environment. A long history of bioinspired algorithms address this problem by crafting rules for navigation that mimic animal behavior: but can effective navigation be learned from the environment rather than set a priori? To address this question we train a reinforcement learning algorithm with realistic turbulent stimuli. Searchers learn to navigate by trial and error and respond solely to odor, with no further input. All computations are defined explicitly, enhancing interpretability. The upshot is that the algorithm identifies odor features as averages over a temporal scale (memory) dictated by the time between odor detections and thus by physics. There is no need to know physics beforehand, as memory can be adjusted based on experience. This approach naturally complements previous algorithms that use prior information and a map of space to plan navigation, rather than learn it from the environment. To what extent different animals plan vs learn to navigate remains to be understood.
Results
@@
In our problem however, the agent does not know where the source is, hence its position $s$ relative to the source, is unknown or “partially observed”. Instead, it can sense odor released by the target. In the language of RL, odor is an “observation” – but does it hold information about the position $s$? The answer is yes: several properties of odor stimuli depend on the distance from the source. However in the presence of turbulence, information lies in the statistics of the odor stimulus. Indeed, when odor is carried by a turbulent flow, it develops into a dramatically stochastic plume, i.e. a complex and convoluted region of space where the fluid is rich in odor molecules. Turbulent plumes break into structures that distort and expand while they travel away from their source and become more and more diluted [31, 4, 32, 6], see Figure 1A. As a consequence, an agent within the plume experiences intermittent odor traces that endlessly switch on (whiff) and off (blank) Figure 1B. The intensity of odor whiffs and how they are interleaved with blanks depends on distance from release, as dictated by physics [32]. Thus the upshot of turbulent transport is that the statistical properties of odor traces depend intricately on the position of the agent relative to the source. In other words, information about the state $s$ is hidden within the observed odor traces.
This positional information can be leveraged with a Bayesian approach that relies on guessing $s$, i.e. defining the probability distribution of the position, also called belief. This is the approach that has been more commonly adopted in the literature until now [26, 27, 28]. Note that because of the complexity of these algorithms, only relatively simple measures of the odor are computationally feasible, e.g. instantaneous presence/absence. Here we take a different model-free and map-free approach. Instead of guessing the current state $s$, we ignore the spatial position and respond directly to the temporal traces of the odor cues. Two other algorithms have been proposed to solve partial observability by responding solely to odor traces with recurrent neural networks [29] and finite state controllers [30] that manipulate implicitly the odor traces. Here instead we manipulate odor traces explicitly, by defining memory as a moving window and by crafting a small number of features of odor traces.
@@
To learn a response to odor traces, we first define a finite set of olfactory states, $o\in O$, so that they bear information about the location $s$. Defining the olfactory states is a challenge due to the dramatic fluctuations and irregularity of turbulent odor traces. To construct a fully interpretable low dimensional state space, we aim at a small number of olfactory states that bear robust information about $s$, i.e. for all values of $s$. We previously found that pairing features of sparsity as well as intensity of turbulent odor traces predicts robustly the location of the source for all $s$ [33]. Guided by these results, we use these two features extracted from the temporal history of odor detections to define a small set of olfactory states.
We proceed to define a function that takes as input the history of odor detections along an agent’s path and returns its current olfactory state. We indicate with $s(t)$ the (unknown) path of an agent, and with $z(s(t),t)$ the observations i.e. odor detections along its path. First, we define a sensing memory $T$ and we consider a short excerpt of the history of odor detections $z_{T}$ of duration $T$ prior to the current time $t$. Formally, $z_{T}(t):=\{z(s(t^{\prime}),t^{\prime})\,|\,t-T\leq t^{\prime}\leq t\}$. Second, we measure the average intensity $c$ (moving average of odor intensity over the time window $T$, conditioned to times when odor is above threshold), and intermittency $i$ (the fraction of time the odor is above threshold during the sensing window $T$). Both features $c$ and $i$ are described by continuous, positive real numbers. Third, we define 15 olfactory states by discretizing $i$ and $c$ in 3 and 5 bins respectively. Intermittency $i$ is bounded between 00 and $1$ and we discretize it in 3 bins by defining two thresholds (33% and 66%). The average concentration, $c$, is bounded between 0 and the odor concentration at the source, hence prior information on the source is needed to discretize $c$ using set thresholds. To avoid relying on prior information, we define thresholds of intensity as percentiles, based on a histogram that is populated online, along each agent’s path (see Materials and Methods). The special case where no odor is detected over $T$ deserves attention, hence we include it as an additional state named “void state” and indicate it with $o\equiv\emptyset$. When $T$ is sufficiently long, the resulting olfactory states map to different spatial locations (Figure 1C, with $T$ equal to the simulation time). Hence this definition of olfactory states can potentially mitigate the problem of partial observability using temporal traces, rather than spatial maps. But will these olfactory states with finite memory $T$ guide agents to the source?
@@
To answer this question, we trained tabular episodic Q learning [18]. In a nutshell, we use a simulator to place an agent at a random location in space at the beginning of each episode. The agent is not aware of its location in space, but it senses odor provided by the fluid dynamics simulator and thus can compute its olfactory state $o$, based on odor detected along its path in the previous $T$ sensing window. It then makes a move according to a set policy of actions $a\sim\pi_{0}(a|o)$. After the move, the simulator displaces the agent to its new location and relays the agent a negative reward $R=-\eta$ if it is not at the source and a positive reward $R=1$ if it reaches the source. The goal of RL is to find a policy of actions that maximizes the expected cumulative future reward $G=E_{\pi}(\sum_{t=0}^{\infty}\gamma^{t}R_{t+1})$ where the expectation is over the ensemble of trajectories and rewards generated by the policy from any initial condition. Because reward is only positive at the source, the optimal policy is the one that reaches the source as fast as possible. To further encourage the agent to reach the source quickly, we introduce a discount factor $\gamma<1$.
@@
Recovery strategy. To fully describe the behavior of our Q-learning agents, we have to prescribe their policy from the void state $o\equiv\emptyset$. This is problematic because turbulent plumes are full of holes thus the void state can occur anywhere both within and outside the plume, Figure 1A. As a consequence, the optimal action $a^{*}(\emptyset)$ from the void state is ill defined. We address this issue by using a separate policy called “recovery strategy”. Inspired by path integration as defined in biology [34, 35, 36], we propose the backtracking strategy consisting in retracing the last $T_{a}$ steps after the agent lost track of the odor. If at the end of backtracking the agent is still in the void state, it activates Brownian motion. Backtracking requires that we introduce memory of the past $T_{a}$ actions. This timescale $T_{a}$ for activating recovery is conceptually distinct from the duration of the sensing memory – however here we set $T_{a}=T$ for simplicity.
@@
We find that Q-learning agents successfully learn to navigate to the odor source by responding solely to their olfactory state, with no sense of space nor models of the odor cues. Learning can be quantified by monitoring the cumulative reward which continuously improves with further training episodes (Figure 1D, left). Improved reward corresponds to agents learning how to reach the source more quickly and reliably with training. Indeed, it is easy to show that the expected cumulative reward $G=\langle e^{-\lambda\tau}-\eta(1-e^{-\lambda\tau})/(1-\gamma)\rangle$, where $\tau$ is a random variable corresponding to time to reach the source and $\gamma=e^{-\lambda\Delta t}$ is the discount factor, with the time step $\Delta t=1$ (see Materials and Methods). Large rewards arise when (i) a large fraction $f^{+}$ of agents successfully reaches the source and (ii) the agents reach the source quickly, which maximizes $g^{+}=\langle e^{-\lambda\tau}|\text{success}\rangle$. Indeed $G=f^{+}G^{+}+(1-f^{+})G^{-}$, where $G^{+}=g^{+}-\eta(1-g^{+})/(1-\gamma)$ and $G^{-}=-\eta(1-e^{-\lambda H_{\text{max}}})/(1-\gamma)$. $H_{\text{max}}$ is the horizon of the agent i.e. the maximum time the agent is allowed to search, and after which the search is considered failed. Note that agents starting closer to the target receive larger rewards purely because of their initial position. To monitor performance independently on the starting location, we introduce the inverse time to reach the source relative to the shortest-path time from the same initial location, which goes from 0 for failing agents to 1 for ideal agents $\langle\tau_{\text{min}}/\tau\rangle$, independently on their starting location. Note that this is not the quantity that is optimized for. One may specifically target this perfomance metrics, which is agnostic on the duration of an agent’s path, by discounting proportionally to $t/\tau_{\text{min}}$.
@@
All four measures of performance plateau to a maximum, suggesting learning has achieved a nearly optimal policy (Figure 1D). Once training is completed, we simulate the trajectory of test-agents starting from any of the about 43 000 admissible locations within the plume and moving according to the optimal policy. We will recapitulate performance with the cumulative reward $G$ averaged over the test-agents and dissect it into speed $g^{+}$, convergence $f^{+}$ and relative time $\langle\tau_{\text{min}}/\tau\rangle$.
Figure 3: The adaptive memory approximates the duration of the blank dictated by physics and it is an efficient heuristics, especially when coupled with a learned recovery strategy. (A) Colormaps of the Eulerian blank times $\tau_{b}$ (top) and the sensing memory $T$ (bottom): Left: averages; Right: standard deviations. The sensing memory statistics is computed over all agents that are located at each discrete cell, at any point in time. (B) Probability distribution of $\tau_{b}$ across all spatial locations and times (black) and of $T$ across all agents at all times (gray). (C) Performance with the adaptive memory nears performance of the optimal fixed memory, here shown for backtracking; similar results apply to the Brownian recovery (Supplementary Figure 2). (D) Comparison of three recovery strategies with adaptive memory: The learned recovery with adaptive memory outperforms all fixed and adaptive memory agents. In (C) and (D) dark squares mark the mean, and light rectangles mark $\pm$ standard deviation. No standard deviation is shown for the $f^{+}$ measure in the learned case as this strategy is deterministic (see Materials and Methods).

### v2

# Q-learning with temporal memory to navigate turbulence

@@ Abstract
We consider the problem of olfactory searches in a turbulent environment. We focus on agents that respond solely to odor stimuli, with no access to spatial perception nor prior information about the odor. We ask whether navigation to a target can be learned robustly within a sequential decision making framework. We develop a reinforcement learning algorithm using a small set of interpretable olfactory states and train it with realistic turbulent odor cues. By introducing a temporal memory, we demonstrate that two salient features of odor traces, discretized in few olfactory states, are sufficient to learn navigation in a realistic odor plume. Performance is dictated by the sparse nature of turbulent odors. An optimal memory exists which ignores blanks within the plume and activates a recovery strategy outside the plume. We obtain the best performance by letting agents learn their recovery strategy and show that it is mostly casting cross wind, similar to behavior observed in flying insects. The optimal strategy is robust to substantial changes in the odor plumes, suggesting minor parameter tuning may be sufficient to adapt to different environments.
1 Introduction
@@
Bacterial cells localize the source of an attractive chemical even if they hold no spatial perception. They respond solely to temporal changes in chemical concentration and the result of their response is that they move toward attractive stimuli by climbing concentration gradients Berg (1975). Larger organisms also routinely sense chemicals in their environment to localize or escape targets, but cannot follow chemical gradients since turbulence breaks odors into sparse pockets and gradients lose information Murlis et al. (1992); Vergassola et al. (2007); Shraiman and Siggia (2000); Balkovsky and I. (2002); Reddy et al. (2022). The question of which features of turbulent odor traces are used by animals for navigation is natural, but not well understood. Beyond olfaction, some animals could use also prior spatial information to navigate  Cardé (2021); Schal (1982); Gire et al. (2016); Baker et al. (2018), but if and how chemosensation and spatial perception are coupled is still not clear.
@@
An algorithmic perspective to olfactory navigation in turbulence can shed light on some of these questions. Without aiming at an exhaustive taxonomy, see e.g. Celani and Panizon (2024) for a recent review, we recall some approaches relevant to put our contribution in context. One class of methods are biomimetic algorithms, where explicit navigation rules are crafted taking inspiration from animal behavior. An advantage of these methods is interpretability, in the sense that they provide insights into features that effectively achieve turbulent navigation, for example: odor presence/absence Baker (1990); Kramer (1997); Belanger and Willis (1988); Balkovsky and I. (2002); odor slope at onset of detection Atema (1996); number of detections in a given interval of time Michaelis et al. (2020) and the time of odor onset Demir et al. (2020a). On the flip side, in biomimetic algorithms behaviors are hardwired and typically reactive, not relying on any optimality criterion.
@@
A way to tackle this shortcoming is to cast olfactory navigation within a sequential decision making framework Sutton and Barto (1998). In this context, navigation is formalized as a task with a reward for success; by maximizing reward, optimal strategies can be sought to efficiently reach the target. A byproduct is that most algorithmic choices can often be done in a principled way. Within this framework, some approaches make explicit use of spatial information. Bayesian algorithms use a spatial map to guess the target location and use odor to refine this guess or “belief”. A prominent algorithm for olfactory navigation based on the concept of belief is the information-seeking algorithm Vergassola et al. (2007) akin to exploration heuristics widely used in robotics Cassandra et al. (1996); LaValle (2006) (see e.g. Loisy and Eloy (2022); Ishida (2020)). Using Bayesian sequential decision making and the notion of beliefs, navigation can be formalized as a Partially Observable Markov Decision Process (POMDP) Krishnamurthy (2016); Hauskrecht (2000); Shani et al. (2013), that colorblackcan be approximatively solved Rigolli et al. (2022b); Heinonen et al. (2023); Loisy and Heinonen (2023b). POMDP approaches are appealing since beliefs are a sufficient statistics for the entire history of odor detections. However, they are computationally cumbersome. Further, they leave the question open of whether navigation as sequential decision making can be performed using solely olfactory information.
@@
Recently, two algorithms studied navigation as a response to olfactory input alone Singh et al. (2023); Verano et al. (2023). In Singh et al. (2023) artificial neural networks were shown to learn near optimal strategies as a response to odor and instantaneous flow direction, although they were trained on odor cues with limited sparsity, and training with sparse odor cues typical of turbulence remains to be tested. In Verano et al. (2023) an approach based on finite state controllers was proposed. Here, optimization was done assuming fixed known mean flow direction, and using a model-based technique, relying on prior knowledge of the likelihood to detect the odor in space, hence still using spatial information. A different model free optimization could also be considered avoiding spatial information but this latter approach also remains to be tested. More generally, all the above approaches manipulate internally the previous history (memory) of odor detections. In this sense they are less interpretable, since the features of odor traces that drive navigation do not emerge explicitly.
@@
In this paper, we propose a reinforcement learning (RL) approach to navigation in turbulence based on a set of interpretable olfactory features, with no spatial information other than the ability to orient relative to the mean flow, and highlight the role played by memory within this context. More precisely, we learn optimal strategies from data by training tabular Q learning Sutton and Barto (1998) with realistic odor cues obtained from state-of-the-art Direct Numerical Simulations of turbulence. From the odor cues, we define features as moving averages of odor intensity and sparsity: the moving window is the temporal memory and naturally connects to the physics of turbulent odors. States are then obtained discretizing such features. Due to sparsity, agents may detect no odor within the moving window. We show there is an optimal memory minimizing the occurrence of this “void state”. The optimal memory scales with the blank time dictated by turbulence as it emerges from a trade off requiring that: (i) short blanks– typical of turbulent plumes– are ignored by responding to detections further in the past, and (ii) long blanks promptly trigger a recovery strategy to make contact with the plume again. We leverage these observations to tune the memory adaptively, by setting it equal to the previous blank experienced along an agent’s path. With this choice, the algorithm tests successfully in distinct environments, suggesting that tuning can be made robustly to enable generalization. The agent learns to surge upwind in most non-void states and to recover by casting crosswind in the absence of detections. Optimal agents limit encounters with the void state to a narrow band right at the edge of the plume. This suggests that the temporal odor features we considered effectively predict when the agent is exiting the plume and point to an intimate connection between temporal predictions and spatial navigation.
Figure 1: Learning a stimulus-response strategy for turbulent navigation. (A) Representation of the search problem with turbulent odor cues obtained from Direct Numerical Simulations of fluid turbulence (grey scale, odor snapshot from the simulations). The discrete position $s$ is hidden; the odor concentration $z_{T}=z(s(t^{\prime}),t^{\prime})|t-T\leq t^{\prime}\leq t$ is observed along the trajectory $s(t^{\prime})$, where $T$ is the sensing memory. (B) Odor traces from direct numerical simulations at different (fixed) points within the plume. Odor is noisy and sparse, information about the source is hidden in the temporal dynamics. (C) Contour maps of olfactory states with nearly infinite memory ($T=2598$): on average olfactory states map to different locations within the plume and the void state is outside the plume. Intermittency is discretized in three bins defined by two thresholds 66% (red line) and 33% (blue line). Intensity is discretized in 5 bins (dark red shade to white shade) defined by four thresholds (percentiles 99%, 80%, 50%, 25%). (D) Performance of stimulus-response strategies obtained during training, averaged over 500 episodes. We train using realistic turbulent data with memory $T=20$ and backtracking recovery.
2 Results
@@
In our problem however, the agent does not know where the source is, hence its position $s$ relative to the source, is unknown or “partially observed”. Instead, it can sense odor released by the target. In the language of RL, odor is an “observation” – but does it hold information about the position $s$? The answer is yes: several properties of odor stimuli depend on the distance from the source Boie et al. (2018); Ackels et al. (2021b); Nag and van Breugel (2024). However in the presence of turbulence, information lies in the statistics of the odor stimulus. Indeed, when odor is carried by a turbulent flow, it develops into a dramatically stochastic plume, i.e. a complex and convoluted region of space where the fluid is rich in odor molecules. Turbulent plumes break into structures that distort and expand while they travel away from their source and become more and more diluted Falkovich et al. (2001); Shraiman and Siggia (2000); Celani et al. (2014); Reddy et al. (2022), see Figure 1A. As a consequence, an agent within the plume experiences intermittent odor traces that endlessly switch on (whiff) and off (blank) Figure 1B. The intensity of odor whiffs and how they are interleaved with blanks depends on distance from release, as dictated by physics Celani et al. (2014). Thus the upshot of turbulent transport is that the statistical properties of odor traces depend intricately on the position of the agent relative to the source. In other words, information about the state $s$ is hidden within the observed odor traces.
This positional information can be leveraged with a Bayesian approach that relies on guessing $s$, i.e. defining the probability distribution of the position, also called belief. This is the approach that has been more commonly adopted in the literature until now Rigolli et al. (2022b); Heinonen et al. (2023); Loisy and Heinonen (2023b). Note that because of the complexity of these algorithms, only relatively simple measures of the odor are computationally feasible, e.g. instantaneous presence/absence. Here we take a different model-free and map-free approach. Instead of guessing the current state $s$, we ignore the spatial position and respond directly to the temporal traces of the odor cues. Two other algorithms have been proposed to solve partial observability by responding solely to odor traces with recurrent neural networks Singh et al. (2023) and finite state controllers Verano et al. (2023) that manipulate implicitly the odor traces. Here instead we manipulate odor traces explicitly, by defining memory as a moving window and by crafting a small number of features of odor traces.
@@
To learn a response to odor traces, we set out to craft a finite set of olfactory states, $o\in O$, so that they bear information about the location $s$. Defining the olfactory states is a challenge due to the dramatic fluctuations and irregularity of turbulent odor traces. To construct a fully interpretable low dimensional state space, we aim at a small number of olfactory states that bear robust information about $s$, i.e. for all values of $s$. We previously found that pairing features of sparsity as well as intensity of turbulent odor traces predicts robustly the location of the source for all $s$ Rigolli et al. (2022a). Guided by these results, we use these two features extracted from the temporal history of odor detections to define a small set of olfactory states.
We proceed to define a function that takes as input the history of odor detections along an agent’s path and returns its current olfactory state. We indicate with $s(t)$ the (unknown) path of an agent, and with $z(s(t),t)$ the observations i.e. odor concentration along its path. First, we define a sensing memory $T$ and we consider a short excerpt of the history of odor detections $z_{T}$ of duration $T$ prior to the current time $t$. Formally, $z_{T}(t):=\{z(s(t^{\prime}),t^{\prime})\,|\,t-T\leq t^{\prime}\leq t\}$. Second, we measure the average intensity $c$ (moving average of odor intensity over the time window $T$, conditioned to times when odor is above threshold), and intermittency $i$ (the fraction of time the odor is above threshold during the sensing window $T$). Both features $c$ and $i$ are described by continuous, positive real numbers. Third, we define 15 olfactory states by discretizing $i$ and $c$ in 3 and 5 bins respectively. Intermittency $i$ is bounded between 00 and $1$ and we discretize it in 3 bins by defining two thresholds (33% and 66%). The average concentration, $c$, is bounded between 0 and the odor concentration at the source, hence prior information on the source is needed to discretize $c$ using set thresholds. To avoid relying on prior information, we define thresholds of intensity as percentiles, based on a histogram that is populated online, along each agent’s path (see Materials and Methods). The special case where no odor is detected over $T$ deserves attention, hence we include it as an additional state named “void state” and indicate it with $o\equiv\emptyset$. When $T$ is sufficiently long, the resulting olfactory states map to different spatial locations (Figure 1C, with $T$ equal to the simulation time). Hence this definition of olfactory states can potentially mitigate the problem of partial observability using temporal traces, rather than spatial maps. But will these olfactory states with finite memory $T$ guide agents to the source?
@@
To answer this question, we trained tabular episodic Q learning Sutton and Barto (1998). In a nutshell, we use a simulator to place an agent at a random location in space at the beginning of each episode. The agent is not aware of its location in space, but it senses odor provided by the fluid dynamics simulator and thus can compute its olfactory state $o$, based on odor detected along its path in the previous $T$ sensing window. It then makes a move according to a set policy of actions $a\sim\pi_{0}(a|o)$. After the move, the simulator displaces the agent to its new location and relays the agent a penalty $R=-\sigma$ with $\sigma=0.001$ if it is not at the source and a reward $R=1$ if it reaches the source. The goal of RL is to find a policy of actions that maximizes the expected cumulative future reward $G=E_{\pi}(\sum_{t=0}^{\infty}\gamma^{t}R_{t+1})$ where the expectation is over the ensemble of trajectories and rewards generated by the policy from any initial condition. Because reward is only positive at the source, the optimal policy is the one that reaches the source as fast as possible. To further encourage the agent to reach the source quickly, we introduce a discount factor $\gamma<1$.
@@
Recovery strategy. To fully describe the behavior of our Q-learning agents, we have to prescribe their policy from the void state $o\equiv\emptyset$. This is problematic because turbulent plumes are full of holes thus the void state can occur anywhere both within and outside the plume, Figure 1A. As a consequence, the optimal action $a^{*}(\emptyset)$ from the void state is ill defined. We address this issue by using a separate policy called “recovery strategy”. Inspired by path integration as defined in biology Etienne and Jeffery (2004); Etienne et al. (1996); Heinze et al. (2018), we propose the backtracking strategy consisting in retracing the last $T_{a}$ steps after the agent lost track of the odor. If at the end of backtracking the agent is still in the void state, it activates Brownian motion. Backtracking requires that we introduce memory of the past $T_{a}$ actions. This timescale $T_{a}$ for activating recovery is conceptually distinct from the duration of the sensing memory – however here we set $T_{a}=T$ for simplicity. Backtracking was observed in ants displaced in unfamiliar environments Wystrach et al. (2013), tsetse flies executing reverse turns bringing them back towards the location where they last detected odor Torr (1988); Gibson and Brady (1985) and cockroaches retracing their steps downwind, sometimes walking all the way back to the release point upon plume loss Willis et al. (2008); it was also previously used in computational models Park et al. (2016).
@@
We find that Q-learning agents successfully learn to navigate to the odor source by responding solely to their olfactory state, with no sense of space nor models of the odor cues. Learning can be quantified by monitoring the cumulative reward which continuously improves with further training episodes (Figure 1D, left). Improved reward corresponds to agents learning how to reach the source more quickly and reliably with training. Indeed, it is easy to show that the expected cumulative reward $G=\langle e^{-\lambda\tau}-\sigma(1-e^{-\lambda\tau})/(1-\gamma)\rangle$, where $\tau$ is a random variable corresponding to time to reach the source and $\gamma=e^{-\lambda\Delta t}$ is the discount factor, with the time step $\Delta t=1$ (see Materials and Methods). Large rewards arise when (i) a large fraction $f^{+}$ of agents successfully reaches the source and (ii) the agents reach the source quickly, which maximizes $g^{+}=\langle e^{-\lambda\tau}|\text{success}\rangle$. Indeed $G=f^{+}G^{+}+(1-f^{+})G^{-}$, where $G^{+}=g^{+}-\sigma(1-g^{+})/(1-\gamma)$ and $G^{-}=-\sigma(1-e^{-\lambda H_{\text{max}}})/(1-\gamma)$. $H_{\text{max}}$ is the horizon of the agent i.e. the maximum time the agent is allowed to search, and after which the search is considered failed. Note that agents starting closer to the target receive larger rewards purely because of their initial position. To monitor performance independently on the starting location, we introduce the inverse time to reach the source relative to the shortest-path time from the same initial location, which goes from 0 for failing agents to 1 for ideal agents $\langle\tau_{\text{min}}/\tau\rangle$, independently on their starting location. Note that this is not the quantity that is optimized for. One may specifically target this perfomance metrics, which is agnostic on the duration of an agent’s path, by discounting proportionally to $t/\tau_{\text{min}}$.
@@
All four measures of performance plateau to a maximum, suggesting learning has achieved a nearly optimal policy (Figure 1D). Once training is completed, we simulate the trajectory of test-agents starting from any of the about 43 000 admissible locations within the plume and moving according to the optimal policy. Admissible locations are defined as any location where the odor is non-zero at least once within the entire simulation. We will recapitulate performance with the cumulative reward $G$ averaged over the test-agents and dissect it into speed $g^{+}$, convergence $f^{+}$ and relative time $\langle\tau_{\text{min}}/\tau\rangle$.